What are three consecutive even integers if the sum of the first, twice the second, and three times the third is equal to 640?
Answer by John10(297)

Let x be the first even integer then the next two consecutive even integers are (x + 2), (x + 4)
Translate to the expression:
x + 2(x + 2) + 3(x + 4) = 640
x + 2x + 4 + 3x + 12 = 640
6x + 16 = 640
6x = 624
x = 104
The the next integers are 106, 108.
john10:)